THE PURE ROTATIONAL SPECTRUM OF TiCl^{+} (X^{3}Φ_{r}) BY VELOCITY MODULATION SPEC-TROSCOPY

\underline{D. T. HALFEN} and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, and Steward Observatory, University of Arizona, Tucson, AZ, 85721.

The pure rotational spectrum of the molecular ion TiCl^{+} (X^{3}Φ_{r}) has been measured using millimeter-wave direct absorption methods incorporating velocity modulation techniques. This species is the first metal-containing molecular ion observed with millimeter-wave velocity modulation spectroscopy. TiCl^{+} was created in an AC glow discharge of gas-phase TiCl_{4} and argon. Ten, eleven, and nine rotational transitions of ^{48}Ti^{35}Cl^{+}, ^{48}Ti^{37}Cl^{+}, and ^{46}Ti^{35}Cl^{+} were measured, respectively, in the frequency range of 323 to 424 GHz. All three spin-orbit components were observed. The irregular fine structure splittings indicate that two spin-orbit ladders are perturbed by an excited ^{3}Δ_{r} electronic state. Rotational, spin-orbit, and spin-spin parameters have been determined from the data and agree well with past optical studies. This study illustrates the power of velocity modulation for ion selectivity at millimeter/sub-mm wavelengths.